Two numbers are in the ratio 2:3. If the sum of their squares is 2925, find these numbers

Answer by Boreal(15235)

4x^2+9x^2=2925
13x^2=2925
x^2=225
x=15 or -15
2x is 30 or -30
3x is 45 or -45
30^2=900
45^2=2025
They add to 2925
Also,(-300^2=900
and (-45^2)=2025
so they also are solutions
30:35
and -30:-45